OpenStudy (anonymous):

Mr. Smith went to visit his daughter. The trip took 7 hours. On the return trip he went by a scenic route which was 20 minutes longer. He drove 5 mph slower and it took him 8 hours. How far did Mr. Smith drive?

OpenStudy (jamesj):

Let x1 be the distance and v1 the first speed. Then 7v1 = x1. On the return trip the distance is x2 = x1 + 20 and v2 = v1 - 5. Then 8v2 = x2 => 8(v1-5) = x1 + 20 You now have two equations in the unknowns v1 and x1. Solve them. Let me know if you get stuck.

OpenStudy (anonymous):

I found 860 from that

OpenStudy (anonymous):

Everything makes sense except the x2 = x1 + 20. Doesn't the 20 need to be multiplied by some rate to make it a distance?

OpenStudy (jamesj):

Oh, sorry yes. x2 = x1 + (20/60)v2 , the /60 is to convert it into hours

OpenStudy (anonymous):

ah, thank you very much! That helped me a lot

OpenStudy (jamesj):

So x2 = x1 + (1/3)(v1 - 5)